OpenStudy (anonymous):

What is the quadratic function that is created with roots -3 and 1 and a vertex at (-1, -8)?

OpenStudy (anonymous):

srry roots r the x values rite

OpenStudy (anonymous):

so we would model this in vertex form

OpenStudy (anonymous):

lets go over this together

OpenStudy (anonymous):

ill wait till u reply

OpenStudy (anonymous):

r u there

OpenStudy (anonymous):

tsunamiiii reply

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so do u remember vertex form.

OpenStudy (anonymous):

what wud the equation be

OpenStudy (anonymous):

f (x) = a(x - h)2 + k ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i mean.. y equals a brackets x minus h... etc

OpenStudy (anonymous):

so it called vertex form for a reason.. (h,k) is the coordinate of the vertex

OpenStudy (anonymous):

so youd insert that in the equation . y = a(x+1)2 - 8

OpenStudy (anonymous):

u can insert a coordinate into the equation and solve for a

OpenStudy (anonymous):

because u already have the (-3,0) and (1,0) , u can sub that into the equation i wrote and solve for the a value

OpenStudy (anonymous):

so try this: 0 = a(1+1)2 - 8

OpenStudy (anonymous):

do u understnad so far

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok tell me what u get as your answer

OpenStudy (anonymous):

u got it..